A survey was given to 140 students. On average, 2 out of 7 students floss every day. How many students do NOT floss every day?


1
Expert's answer
2021-05-07T12:05:05-0400

If on average 2 out of 7 students use floss every day, then among 140 students use floss every day

"140 \cdot \frac{2}{7} = 40" students.

Then don't floss every day 140-40=100 students.

Answer: 100 students.